Select the correct graph of the given function, which is assumed to be zero outside the given interval. Then find its transfoFind the transform and sketch the correct graph clearly I will upvote
Answer #1

By transform if you mean Laplace transform then

Given function is f(t)=\cos(8t)~~~~,~~~0<t<\pi

By definition of Laplace transform we get

\frak{L} (f(t))=\int_{0}^{ \infty} f(t) e^{-st} dt \\ \\~~~ ~~~~~~~~~~~~~= \int_{0}^{\pi}\cos (8t)e^{-st }dt\\~~~ ~~~~~~~~~~~~~= \left [ {e^{-st}\over 8^2+(-s)^2}[8\sin(8t)+(-s)\cos (8t)] \right ]_{0}^{\pi} \\ \\~~~ ~~~~~~~~~~~~~={s[1-e^{-\pi s}]\over s^2+64}